MILLIMETER-WAVE SPECTRUM OF ISO-PROPANOL

\underline{A. MAEDA}, F. C. DE LUCIA, E. HERBST, Department of Physics, The Ohio State University, Columbus, OH 43210.

With the aim of astronomical observation, the millimeter-wave spectrum of iso-propanol has been observed in the 100-370 GHz region with the FASSST spectrometer. Over 11000 transitions within three torsional substates and between two gauche substates have been successfully assigned up to J′′ = 68 with the CAAARS program. The rotation of the hydroxyl portion gives iso-propanol three stable configurations labeled trans, gauche and gauche’ corresponding to the dihedral angle of H-C-O-H. The two equivalent gauche forms interact through tunneling to split the degeneracy in energy. The resulting torsional substates can be characterized by the labels gs (symmetric gauche) and ga (anti-symmetric gauche). Previous microwave and infrared spectroscopic studies concluded that gs is the global minimum and trans is the highest in energy. In our new simultaneous fit of gs and ga data using a two-state Hamiltonian, it was necessary to include up to fifth-order torsion-rotation interaction terms as well as sextic centrifugal distortion terms to analyze the millimeter-wave spectrum of iso-propanol. The lease-square fit of gauche resulted in a RMS of 88 kHz. The trans substate was analyzed with a rotational Hamiltonian for a quasi-rigid molecule with up to sextic centrifugal distortion terms with a RMS of the fit 63 kHz. Despite a number of theoretical studies as well as experimental work, the torsional energy of the trans conformer has not yet been determined because of a lack of direct measurements of torsional transitions between the trans and gauche substates. We estimated the trans-gauche energy separation from relative intensity data.